Show that 𝑑𝑦/𝑑π‘₯ = 𝑠𝑒𝑐^2 x given that 𝑦 = tan x


1
Expert's answer
2021-06-07T16:21:14-0400
"\dfrac{d}{dx}(\tan (x))=\lim\limits_{h\to0}\dfrac{\tan(x+h)-\tan(x)}{h}"


"=\lim\limits_{h\to0}\dfrac{\sin(x+h)\cos(x)-\cos(x+h)\sin(x)}{h \cos(x+h)\cos(x)}"

"=\lim\limits_{h\to0}\dfrac{\sin(x+h-x)}{h \cos(x+h)\cos(x)}"


"=\lim\limits_{h\to0}\dfrac{\sin(x+h-x)}{h \cos(x+h)\cos(x)}"

"=\lim\limits_{h\to0}\dfrac{\sin(h)}{h }\cdot\lim\limits_{h\to0}\dfrac{1}{ \cos(x+h)\cos(x)}"

"=1\cdot\lim\limits_{h\to0}\dfrac{1}{h \cos(x+h)\cos(x)}"


"=\dfrac{1}{ \cos(x+0)\cos(x)}"

"=\sec^2 (x)"

"\dfrac{d}{dx}(\tan (x))=\sec^2(x)"